Suppose your firm produces electricity by burning coal. currently it buys Central Appalachia 12500 BTU per ton coal at the market price of $52 per ton. Average cost to produce coal is 35 per ton. So your firm considers buying the mine and saving $17 per ton of coal burned. Each power plant burns 1.4 million tons of coal per year.

Is this a good idea? Why or why not?
Answer #1

Answer : Yes, this is a good idea.

Because here the market price of coal is $52 per ton of coal. Each power plant requires 1.4 million or 1,400,000 tons of coal. So, at market price the per year cost for each power plant's electricity production = 52 * 1,400,000 = $72,800,000 = $72.8 million.

Now if my firm buy the mine then the cost for per ton coal production is $35. So, the per year cost for each power plant's electricity production = 35 * 1,400,000 = $49,000,000 = $49 million.

So, if the firm buy the mine then the firm can save (72.8 - 49) = $23.8 million per year.

As here by purchasing the mine the firm can save a massive amount per year hence purchasing the mine is a good idea for my firm.